Find p^ and q^ X = 28 n = 110


1
Expert's answer
2021-05-23T23:19:04-0400

"p=\dfrac{X}{n}=\dfrac{28}{110}=\dfrac{14}{55}"

"q=1-p=1-\dfrac{14}{55}=\dfrac{41}{55}"